# Neural Configuration-Space Barriers for Manipulation Planning and Control

### v1

@@ I Introduction
  • We develop a motion planning algorithm that uses a neural CDF barrier to construct a graph of fully traversable spheres in the robot’s configuration space, allowing significantly reduced collision checks.
  • We purpose a distributionally robust control barrier function formulation from the neural CDF barrier to account for dynamic obstacles, CDF modeling inaccuracies, and sensor noise.
  • We validate our CDF-based planning and control algorithms in extensive simulations and real-robot experiments on a 6-DoF xArm manipulator in dynamic and cluttered environments.
@@ II Related Work
III Notation
The sets of real, non-negative real, and natural numbers are denoted by $\mathbb{R}$, $\mathbb{R}_{\geq 0}$, and $\mathbb{N}$, respectively. For $N\in\mathbb{N}$, we let $[N]:=\{1,2,\dots N\}$. The interior and boundary of a set $\mathcal{C}\subset\mathbb{R}^{n}$ are denoted by $\text{Int}(\mathcal{C})$ and $\partial\mathcal{C}$. The gradient of a differentiable function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is denoted by $\nabla V$. A continuous function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ is of extended class $\mathcal{K}_{\infty}$ if it is strictly increasing, $\alpha(0)=0$, and $\lim_{r\rightarrow\infty}\alpha(r)=\infty$. The special orthogonal group of dimension $p$ is denoted by $\text{SO}(p)=\{\mathbf{R}\in\mathbb{R}^{p\times p}\mid\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}_{p},\det(\mathbf{R})=1\}$. The special Euclidean group of dimension $p$, denoted $\text{SE}(p)$ is the set of matrices $\begin{pmatrix}\mathbf{R}&\mathbf{t}\\ \mathbf{0}^{\top}&1\end{pmatrix}$ with $\mathbf{R}\in\text{SO}(p)$ and $\mathbf{t}\in\mathbb{R}^{p}$.
IV Problem Formulation
We consider a robot manipulator with $m$ revolute joints and $m+1$ links. The joint configuration is represented by $\mathbf{q}=[q_{1},\ldots,q_{m}]^{\top}\in\mathcal{Q}$, where $\mathcal{Q}=\{\mathbf{q}\in\mathbb{R}^{m}\mid\mathbf{q}_{\min}\leq\mathbf{q}\leq\mathbf{q}_{\max}\}$ is the configuration space bounded by the joint limits. The forward kinematics of the manipulator map a joint configuration $\mathbf{q}$ to the end-effector pose in task space, represented as $\mathbf{T}_{\text{ee}}(\mathbf{q})\in\text{SE}(3)$.
Given a configuration $\mathbf{q}$, we denote the robot shape by a set-valued function $\mathcal{S}(\mathbf{q})\subset\mathbb{R}^{3}$, and its surface by $\partial\mathcal{S}(\mathbf{q})$. The kinematics of the manipulator are described by:
@@ IV Problem Formulation
We consider a dynamic environment containing both static and dynamic obstacles. At time $t$, let $\mathcal{O}(t)\subset\mathbb{R}^{3}$ denote the closed set of obstacles, and let the free space be represented as $\mathcal{F}(t)=\mathbb{R}^{3}\setminus\mathcal{O}(t)$, which is an open set. We assume that a sensor (e.g., depth camera or LiDAR) provides point-cloud measurements $\mathcal{P}(t)\subset\mathbb{R}^{3}$, which approximate the boundary of the time-varying obstacle region $\mathcal{O}(t)$. The extrinsic parameters relating the sensor pose to the robot base are assumed to be calibrated offline.
@@
Consider a robot manipulator with initial configuration $\mathbf{q}_{0}$ and dynamics in (1). Given point-cloud observations $\mathcal{P}(t)$ of the obstacles $\mathcal{O}(t)$ in the environment, design a control policy that drives the robot’s end-effector $\mathbf{T}_{\text{ee}}$ to a desired goal pose $\mathbf{T}_{\text{G}}\in\text{SE}(3)$, while ensuring the robot body remains in free space, i.e., $\mathcal{S}(\mathbf{q}(t))\subset\mathcal{F}(t)$, $\forall t\geq 0$.
V Safe Barrier from Configuration-Space Distance Function
Efficient motion planning and control for robotic manipulators require accurate and scalable safety representations in configuration space. In this section, we first review the notion of a signed distance function (SDF) in the workspace, then introduce the CDF and its key properties, and finally define a neural CDF barrier for planning and control applications.
V-A Configuration Space Distance Function
Motivated by successful applications in safe mobile robot navigation [17, 30, 22], SDFs have recently been used for motion planning and control of robot manipulators [23, 28, 45, 31, 32]. An SDF $f_{s}$ measures the distance from a workspace point $\mathbf{p}\in\mathbb{R}^{3}$ to the robot surface $\partial\mathcal{S}(\mathbf{q})$:
$f_{s}(\mathbf{p},\mathbf{q})=\begin{cases}-\min\limits_{\mathbf{p}^{*}\in\partial\mathcal{S}(\mathbf{q})}\|\mathbf{p}-\mathbf{p}^{*}\|,&\text{if }\mathbf{p}\in\mathcal{S}(\mathbf{q}),\\ \phantom{-}\min\limits_{\mathbf{p}^{*}\in\partial\mathcal{S}(\mathbf{q})}\|\mathbf{p}-\mathbf{p}^{*}\|,&\text{if }\mathbf{p}\notin\mathcal{S}(\mathbf{q}).\end{cases}$ (2)
@@ V-A Configuration Space Distance Function
By operating directly in joint space, CDF-based methods avoid iterative conversions between task and configuration spaces, and provide gradient information that is inherently meaningful for optimization-based joint control synthesis [7].
V-B Neural CDF Barrier
Building on the concept of the CDF, we define a CDF barrier to ensure the robot remains in a collision-free region of the configuration space. Formally, a CDF barrier associated with robot CDF $f_{c}$ and obstacle set $\mathcal{O}(t)$ is defined as:
$h(\mathbf{q},t):=\inf_{\mathbf{p}\in\partial\mathcal{O}(t)}f_{c}(\mathbf{p},\mathbf{q}),$ (5)
where $\partial\mathcal{O}(t)$ is the obstacle set boundary. Then, the time-varying safe set in configuration space induced by $h(\mathbf{q},t)$ is
@@ V-B Neural CDF Barrier
In real-world settings, $\partial\mathcal{O}(t)$ must be inferred from point-cloud measurements $\mathcal{P}(t)$ from a depth camera or LiDAR. Because the points in $\mathcal{P}(t)$ are noisy samples from the boundary of $\mathcal{O}(t)$, we approximate the CDF barrier $h(\mathbf{q},t)$ as the minimum of $f_{c}(\mathbf{p},\mathbf{q})$ over the points in $\mathcal{P}(t)$:
$h(\mathbf{q},t)\;\approx\;\min_{\mathbf{p}\,\in\,\mathcal{P}(t)}f_{c}(\mathbf{p},\mathbf{q}).$ (7)
Moreover, computing $f_{c}(\mathbf{p},\mathbf{q})$ exactly can be difficult for high-DoF manipulators because it involves (infinitely) many potential contact configurations $\mathbf{q}^{\prime}$ in (3). Following Li et al. [27], we model $f_{c}$ by a multi-layer perceptron $\hat{f}_{c}(\mathbf{p},\mathbf{q};\boldsymbol{\theta})$ with learnable parameters $\boldsymbol{\theta}$. This representation reduces the storage requirements compared to volumetric or tabular encodings and supports parallel queries for a set of points $\mathbf{p}$ or configurations $\mathbf{q}$ with associated gradient computations. However, learning-based CDF representations inevitably introduce modeling errors. Combining the point-cloud approximation with the neural CDF approximation yields a practical CDF barrier:
$\hat{h}(\mathbf{q},t;\boldsymbol{\theta})\;=\;\min_{\mathbf{p}\,\in\,\mathcal{P}(t)}\hat{f}_{c}(\mathbf{p},\mathbf{q};\,\boldsymbol{\theta}),$ (8)
where the noise in $\mathcal{P}(t)$, its lack of complete coverage of the obstacle boundary, and the approximation error in $\hat{f}_{c}(\cdot;\boldsymbol{\theta})$ constitute multiple sources of uncertainty.
In the following sections, we leverage the fast parallel query capabilities of the neural CDF barrier to enable efficient sampling-based planning in configuration space (Sec. VI), and then propose a distributionally robust control strategy that tracks the planned trajectory while ensuring real-time safety despite uncertainties and dynamic obstacles (Sec. VII).
VI Configuration Space Bubble-Based Planning
@@
VI-A Rapidly-exploring Bubble Graph (RBG)
Suppose $\mathbf{q}\in\mathcal{Q}$ is a safe configuration. We aim to identify a radius $r(\mathbf{q})\geq 0$ such that any configuration $\mathbf{q}^{\prime}$ satisfying $\|\mathbf{q}-\mathbf{q}^{\prime}\|\leq r(\mathbf{q})$ is safe. We call this spherical region a configuration-space bubble and denote it as $\mathcal{B}(\mathbf{q},r(\mathbf{q}))$. Given the neural CDF barrier $\hat{h}(\mathbf{q},t)$ for fixed time $t$, we construct the bubble radius at configuration $\mathbf{q}$ as:
@@ VI-A Rapidly-exploring Bubble Graph (RBG)
Algorithm 1 details how RBG constructs a bubble-based graph. RBG iteratively samples random configurations (biased toward goal configurations), identifies the nearest bubble in configuration space, and if the sample lies outside that bubble, creates a new bubble by querying the neural CDF barrier to determine its radius. The newly created bubble is then connected to any existing bubbles $\mathcal{B}(\mathbf{q}_{i},r(\mathbf{q}_{i}))$ with overlapping safe regions, i.e., if
@@
VI-B Path Selection and Trajectory Optimization
@@
Thus far, we have focused on global planning in static environments, assuming a time-invariant and accurate neural CDF barrier $\hat{h}(\mathbf{q},t)$. However, real-world deployments must contend with dynamic obstacles, CDF modeling errors, and sensor uncertainties. In Section VII, we introduce a distributionally robust control formulation to track the planned path in real time, explicitly accounting for these uncertainties to ensure safe and adaptive execution.
VII Distributionally Robust Manipulation Control
In this section, we present a control synthesis approach that guarantees safe real-time configuration trajectory tracking using imperfect neural CDF barrier estimates (due to sensing and modeling errors) even in the presence of dynamic obstacles. We first describe our approach to track the reference trajectory (Sec. VII-A). We then incorporate a distributionally robust control barrier function (DR-CBF) [33] to ensure safe control synthesis despite dynamic obstacles and neural CDF barrier uncertainty (Sec. VII-B).
VII-A Governor-based Trajectory Tracking
We first present our approach for tracking the planned piecewise Bézier configuration trajectory $\gamma:[0,1]\to\mathcal{Q}$, obtained in Sec. VI-B, in the absence of safety constraints. Inspired by Garone and Nicotra [16], Li et al. [29], we design a reference governor to determine a dynamically changing reference configuration $\gamma(s)$ along the trajectory, where $s\in[0,1]$ is a progress variable. In other word, $\gamma(0)=\mathbf{q}_{0}$ is the initial robot configuration, while $\gamma(1)=\mathbf{q}_{\text{G}}$ is one of the goal configurations. The governor system adapts the value of the progress variable $s$ based on the manipulator’s current configuration to ensure smooth progress along the trajectory. We then stabilize the manipulator configuration $\mathbf{q}$ to $\gamma(s)$ using a proportional-derivative (PD) controller.

### v2

@@ I Introduction
  • We introduce a unified neural CDF barrier formulation that estimates the configuration-space distance to both environment obstacles and self-collisions. This is achieved by separately learning an environment CDF and a self-collision CDF for high-DoF manipulators. The resulting barrier enables efficient and scalable safety evaluation based on point-cloud observations.
  • We develop the bubble-CDF motion planning algorithm, which constructs a graph of fully traversable configuration-space bubbles using the neural CDF barrier. This approach significantly reduces the number of collision checks and achieves faster planning times compared to traditional sampling-based planners.
  • We propose a distributionally robust control barrier function (DR-CBF) formulation based on the neural CDF barrier, which accounts for dynamic obstacles, modeling errors, and perception uncertainty.
  • We validate the proposed framework in both simulation and real-robot experiments with a 6-DoF xArm manipulator operating in dynamic and cluttered environments, demonstrating improved efficiency and safety.
@@ II Related Work
00footnotetext: Notation: The sets of real, non-negative real, and natural numbers are denoted by $\mathbb{R}$, $\mathbb{R}_{\geq 0}$, and $\mathbb{N}$, respectively. For $N\in\mathbb{N}$, we let $[N]:=\{1,2,\dots N\}$. The interior and boundary of a set ${\cal C}\subset\mathbb{R}^{n}$ are denoted by $\text{Int}({\cal C})$ and $\partial{\cal C}$. The gradient of a differentiable function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is denoted by $\nabla V$. A continuous function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ is of extended class ${\cal K}_{\infty}$ if it is strictly increasing, $\alpha(0)=0$, and $\lim_{r\rightarrow\infty}\alpha(r)=\infty$. The special orthogonal group of dimension $p$ is denoted by $\text{SO}(p)=\{\mathbf{R}\in\mathbb{R}^{p\times p}\mid\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}_{p},\det(\mathbf{R})=1\}$. The special Euclidean group of dimension $p$, denoted $\text{SE}(p)$ is the set of matrices $\begin{pmatrix}\mathbf{R}&\mathbf{t}\\ \mathbf{0}^{\top}&1\end{pmatrix}$ with $\mathbf{R}\in\text{SO}(p),\;\mathbf{t}\in\mathbb{R}^{p}$.
III Problem Formulation
We consider a robot manipulator with $m$ revolute joints and $m+1$ links, where Link 0 denotes the fixed base. The joint configuration is represented by $\mathbf{q}=[q_{1},\ldots,q_{m}]^{\top}\in{\cal Q}$, where ${\cal Q}=\{\mathbf{q}\in\mathbb{R}^{m}\mid\mathbf{q}_{\min}\leq\mathbf{q}\leq\mathbf{q}_{\max}\}$ is the configuration space bounded by the joint limits. The forward kinematics of the manipulator map a joint configuration $\mathbf{q}$ to the end-effector pose in task space, represented as $\mathbf{T}_{\text{ee}}(\mathbf{q})\in\text{SE}(3)$.
Given a configuration $\mathbf{q}$, we denote the robot shape by a set-valued function ${\cal S}(\mathbf{q})\subset\mathbb{R}^{3}$, and its surface by $\partial{\cal S}(\mathbf{q})$. The kinematics of the manipulator are described by:
@@ III Problem Formulation
We consider a dynamic environment containing both static and dynamic obstacles. At time $t$, let $\mathcal{O}(t)\subset\mathbb{R}^{3}$ denote the closed set of obstacles, and let the free space be represented as $\mathcal{F}(t)=\mathbb{R}^{3}\setminus\mathcal{O}(t)$, which is an open set. We assume that a sensor (e.g., depth camera or LiDAR) provides point-cloud measurements ${\cal P}(t)\subset\mathbb{R}^{3}$, which approximate the boundary of the time-varying obstacle region $\mathcal{O}(t)$. The extrinsic parameters relating the sensor pose to the robot base are assumed to be calibrated offline.
@@
Consider a robot manipulator with initial configuration $\mathbf{q}_{0}$ and dynamics in (1). Given point-cloud observations ${\cal P}(t)$ of the obstacles ${\cal O}(t)$ in the environment, design a control policy that drives the robot’s end-effector $\mathbf{T}_{\text{ee}}$ to a desired goal pose $\mathbf{T}_{\text{G}}\in\text{SE}(3)$, while ensuring the robot body remains in free space, i.e., $\mathcal{S}(\mathbf{q}(t))\subset\mathcal{F}(t)$, $\forall t\geq 0$.
IV Safe Barrier from Configuration-Space Distance Function
Efficient motion planning and control for robotic manipulators require accurate and scalable safety representations in configuration space. In this section, we first review the notion of a signed distance function (SDF) in the workspace, then introduce the environment and self-collision CDF and their key properties, and finally define a neural CDF barrier for planning and control applications.
IV-A Environment Configuration Space Distance Function
Motivated by successful applications in safe mobile robot navigation [17, 32, 24], SDFs have recently been used for motion planning and control of robot manipulators [25, 30, 46, 33]. An SDF $f_{s}$ measures the distance from a workspace point $\mathbf{p}\in\mathbb{R}^{3}$ to the robot surface $\partial{\cal S}(\mathbf{q})$:
$f_{s}(\mathbf{p},\mathbf{q})=\begin{cases}-\min\limits_{\mathbf{p}^{*}\in\partial{\cal S}(\mathbf{q})}\|\mathbf{p}-\mathbf{p}^{*}\|,&\text{if }\mathbf{p}\in{\cal S}(\mathbf{q}),\\ \phantom{-}\min\limits_{\mathbf{p}^{*}\in\partial{\cal S}(\mathbf{q})}\|\mathbf{p}-\mathbf{p}^{*}\|,&\text{if }\mathbf{p}\notin{\cal S}(\mathbf{q}).\end{cases}$ (2)
@@ IV-A Environment Configuration Space Distance Function
In practice, computing (3) requires identifying a set of zero-level configurations $\mathbf{q}^{\prime}$ satisfying $f_{s}(\mathbf{p},\mathbf{q}^{\prime})=0$. These configurations are typically obtained using numerical optimization methods such as Broyden-Fletcher-Goldfarb–Shanno (BFGS) algorithm [20]. However, due to the high dimensionality of the configuration space and the sparsity of $\mathbf{q}^{\prime}$ samples, this direct formulation can lead to an overly smooth or inaccurate approximation of the true CDF. Therefore, Li et al. [29] leverage the fact that the contact at a point $\mathbf{p}$ is primarily determined by a subset of joints preceding the contact link. Let $k$ denote the robot link that comes into contact with $\mathbf{p}$, and let $\mathbf{q}_{:k}$ represent the joint angles influencing the motion of link $k$. The CDF is then refined by restricting the distance computation to these relevant joints:
$f_{c}(\mathbf{p},\mathbf{q})=\min_{k=1,\dots,m}\min_{\mathbf{q}^{\prime}}\|\mathbf{q}_{:k}-\mathbf{q}^{\prime}_{:k}\|,\quad\text{s.t.}\quad f_{s}(\mathbf{p},\mathbf{q}^{\prime})=0,$ (5)
where $m$ is the total number of robot joints.
IV-B Self-Collision Configuration Space Distance Function
While environment CDFs measure proximity to external obstacles, self-collisions between different parts of the robot pose a distinct challenge, especially for high-DoF manipulators where self-intersections are configuration-dependent and complex. To address this, we define the Self-Collision Configuration Distance Function (SCDF), which quantifies the minimum joint-space distance from a configuration $\mathbf{q}$ to the set of self-colliding configurations.
Let $\mathcal{C}_{\mathrm{sc}}\subset{\cal Q}$ denote the closed set of joint configurations that result in self-collision. Then, for a given configuration $\mathbf{q}\in{\cal Q}$, we define the SCDF as:
$f_{\mathrm{sc}}(\mathbf{q}):=\min_{\mathbf{q}^{\prime}\in\mathcal{C}_{\mathrm{sc}}}\|\mathbf{q}-\mathbf{q}^{\prime}\|.$ (6)
Similar to Section IV-A, the SCDF in (6) represents the Euclidean distance from $\mathbf{q}$ to the closed set of self-colliding configurations $\mathcal{C}_{\mathrm{sc}}\subset\mathbb{R}^{m}$. Note that $\mathcal{C}_{\mathrm{sc}}$ may consist of multiple disconnected components, as different joint subsets can lead to distinct types of self-collision. Nevertheless, the SCDF inherits the Eikonal property and satisfies
$\|\nabla_{\mathbf{q}}f_{\mathrm{sc}}(\mathbf{q})\|=1.$ (7)
In practice, $\mathcal{C}_{\mathrm{sc}}$ is often only sparsely sampled, and self-collisions typically involve only a subset of joints. As such, computing the full-joint distance in (6) can lead to overly conservative and inaccurate approximations. To improve accuracy and better exploit the robot’s kinematic structure, we refine the SCDF by computing distance only over the subset of joints responsible for the self-collision.
Specifically, if a self-collision at $\mathbf{q}^{\prime}$ occurs between Link $i$ and Link $j$, we define the joint subvector $\mathbf{q}_{a:b}$, where $a=\min(i,j)$ and $b=\max(i,j)$, to include only joints that affect the relative motion of these links. The refined SCDF becomes:
$f_{\mathrm{sc}}(\mathbf{q}):=\min_{\mathbf{q}^{\prime}\in\mathcal{C}_{\mathrm{sc}}}\|\mathbf{q}_{a:b}-\mathbf{q}^{\prime}_{a:b}\|.$ (8)
For example, if the collision occurs between Link 3 and Link 5, then joints 3 through 5 may primarily determine their relative positions, so the distance is computed using $\mathbf{q}_{3:5}$.
IV-C Neural CDF Barrier
Building on the environment and self-collision CDF, we define a CDF barrier to ensure the robot remains in a collision-free region of the configuration space. Formally, a CDF barrier associated with environment CDF $f_{c}$, self-collision CDF $f_{sc}$, and obstacle set ${\cal O}(t)$ is defined as:
$h(\mathbf{q},t):=\min\{\inf_{\mathbf{p}\in\partial\mathcal{O}(t)}f_{c}(\mathbf{p},\mathbf{q}),f_{sc}(\mathbf{q})\}$ (9)
where $\partial{\cal O}(t)$ is the obstacle set boundary. Then, the time-varying safe set in configuration space induced by $h(\mathbf{q},t)$ is
@@ IV-C Neural CDF Barrier
In real-world settings, $\partial\mathcal{O}(t)$ must be inferred from point-cloud measurements ${\cal P}(t)$ from a depth camera or LiDAR. Because the points in ${\cal P}(t)$ are noisy samples from the boundary of ${\cal O}(t)$, we approximate the CDF barrier as
$h(\mathbf{q},t)\;\approx\;\min\{\min_{\mathbf{p}\,\in\,\mathcal{P}(t)}f_{c}(\mathbf{p},\mathbf{q}),f_{sc}(\mathbf{q})\}.$ (11)
Moreover, computing $f_{c}(\mathbf{p},\mathbf{q})$ and $f_{sc}(\mathbf{q})$ exactly are difficult for high-DoF manipulators because it involves (infinitely) many potential contact configurations $\mathbf{q}^{\prime}$ in (3) and self-collision configurations in (8). Similar to Li et al. [29], we model $f_{c}$ and $f_{sc}$ by multi-layer perceptrons $\hat{f}_{c}(\mathbf{p},\mathbf{q};\boldsymbol{\theta_{1}})$ and $\hat{f}_{sc}(\mathbf{p},\mathbf{q};\boldsymbol{\theta_{2}})$ with learnable parameters $\boldsymbol{\theta_{1}}$ and $\boldsymbol{\theta_{2}}$, respectively. This representation reduces the storage requirements compared to volumetric or tabular encodings and supports parallel queries for a set of points $\mathbf{p}$ or configurations $\mathbf{q}$ with associated gradient computations.
However, learning-based representations inevitably introduce modeling errors. Combining the point-cloud approximation with the neural CDF approximation yields a practical CDF barrier:
$\hat{h}(\mathbf{q},t;\boldsymbol{\theta})\;=\;\min\{\min_{\mathbf{p}\,\in\,\mathcal{P}(t)}\hat{f}_{c}(\mathbf{p},\mathbf{q};\,\boldsymbol{\theta_{1}}),\hat{f}_{sc}(\mathbf{q};\boldsymbol{\theta_{2}})\}$ (12)
where the noise in $\mathcal{P}(t)$, its lack of complete coverage of the obstacle boundary, and the approximation error in $\hat{f}_{c}$ and $\hat{f}_{sc}$ constitute multiple sources of uncertainty.
In the following sections, we leverage the fast parallel query capabilities of the neural CDF barrier to enable efficient sampling-based planning in configuration space (Sec. V), and then propose a distributionally robust control strategy that tracks the planned trajectory while ensuring real-time safety despite uncertainties and dynamic obstacles (Sec. VI).
V Configuration Space Bubble-Based Planning
@@
V-A Rapidly-exploring Bubble Graph (RBG)
Suppose $\mathbf{q}\in{\cal Q}$ is a safe configuration. We aim to identify a radius $r(\mathbf{q})\geq 0$ such that any configuration $\mathbf{q}^{\prime}$ satisfying $\|\mathbf{q}-\mathbf{q}^{\prime}\|\leq r(\mathbf{q})$ is safe. We call this spherical region a configuration-space bubble and denote it as $\mathcal{B}(\mathbf{q},r(\mathbf{q}))$. Given the neural CDF barrier $\hat{h}(\mathbf{q},t)$ for fixed time $t$, we construct the bubble radius at configuration $\mathbf{q}$ as:
@@ V-A Rapidly-exploring Bubble Graph (RBG)
Algorithm 1 details how RBG constructs a bubble-based graph. RBG iteratively samples random configurations (biased toward goal configurations), identifies the nearest bubble in configuration space, and if the sample lies outside that bubble, creates a new bubble by querying the neural CDF barrier to determine its radius. The newly created bubble is then connected to any existing bubbles ${\cal B}(\mathbf{q}_{i},r(\mathbf{q}_{i}))$ with overlapping safe regions, i.e., if
@@
V-B Path Selection and Trajectory Optimization
@@
Thus far, we have focused on global planning in static environments, assuming a time-invariant and accurate neural CDF barrier $\hat{h}(\mathbf{q},t)$. However, real-world deployments must contend with dynamic obstacles, CDF modeling errors, and sensor uncertainties. In Section VI, we introduce a distributionally robust control formulation to track the planned path in real time, explicitly accounting for these uncertainties to ensure safe and adaptive execution.
VI Distributionally Robust Manipulation Control
In this section, we present a control synthesis approach that guarantees safe real-time configuration trajectory tracking using imperfect neural CDF barrier estimates (due to sensing and modeling errors) even in the presence of dynamic obstacles. We first describe our approach to track the reference trajectory (Sec. VI-A). We then incorporate a distributionally robust control barrier function (DR-CBF) [34] to ensure safe control synthesis despite dynamic obstacles and neural CDF barrier uncertainty (Sec. VI-B).
VI-A Governor-based Trajectory Tracking
We first present our approach for tracking the planned piecewise Bézier configuration trajectory $\gamma:[0,1]\to{\cal Q}$, obtained in Sec. V-B, in the absence of safety constraints. Inspired by Garone and Nicotra [16], Li et al. [31], we design a reference governor to determine a dynamically changing reference configuration $\gamma(s)$ along the trajectory, where $s\in[0,1]$ is a progress variable. In other word, $\gamma(0)=\mathbf{q}_{0}$ is the initial robot configuration, while $\gamma(1)=\mathbf{q}_{\text{G}}$ is one of the goal configurations. The governor system adapts the value of the progress variable $s$ based on the manipulator’s current configuration to ensure smooth progress along the trajectory. We then stabilize the manipulator configuration $\mathbf{q}$ to $\gamma(s)$ using a proportional-derivative (PD) controller.